OpenStudy (anonymous):

: Show all work. Find the following quotient. 25x^2 – 100y^2/5x + 10y

OpenStudy (dumbcow):

\[\frac{25x^{2}-100y^{2}}{5x+10y} = \frac{25(x^{2}-4y^{2})}{5(x+2y)} = \frac{5(x+2y)(x-2y)}{x+2y} = 5(x-2y)\]

jimthompson5910 (jim_thompson5910):

\[\frac{25x^2-100y^2}{5x+10y}\] \[\frac{(5x+10y)(5x-10y)}{5x+10y}\] \[5x-10y\] So \[\frac{25x^2-100y^2}{5x+10y}=5x-10y\]

OpenStudy (anonymous):

thank you